40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 8647     Entry time: Tue May 28 11:11:37 2013
Author: Manasa 
Type: Update 
Category: IOO 
Subject: MC aligned 

[Jenne, Manasa]

Fixed crappy alignment of MC by moving MC mirrors. MC REFL PD measured 0.5 after alignment. Spot positions were measured using msassMCdecenter. Plot for the same is attached.

Attachment 1: MC_0528.png  173 kB  | Hide | Hide all
MC_0528.png
ELOG V3.1.3-